# Deep Neural Network-Based Voltage Prediction for Alkali-Metal-Ion Battery Materials

### v1

@@ 3.5 Model Predicted vs DFT Calculated Voltage
3.6 Conclusions
@@
This study demonstrates the power of combining machine learning (ML) with first-principles simulations to accelerate the discovery and optimization of next-generation Na-ion battery cathodes. Our deep neural network (DNN) model, trained on Density Functional Theory (DFT) data, achieves a mean absolute error (MAE) of 0.24V, outperforming existing voltage prediction models with a high level of accuracy and a deviation of less than 5%. To validate the predictive capability of our model, we compared its voltage estimations with DFT-calculated values for two newly investigated cathode compositions: O3-Na0.67Mn0.61Fe0.28Cr0.11O2 and P2-Na0.67Mn0.5Fe0.28Cu0.22O2. The excellent agreement between ML-predicted and DFT-computed voltages confirms the reliability of our approach in capturing the electrochemical behavior of Na-ion materials. By significantly reducing reliance on time-consuming and resource-intensive experimental methods, our approach provides a systematic framework for identifying promising cathode materials with optimized electrochemical properties. This data-driven methodology enables rapid screening of novel compositions, guiding experimental efforts toward the most promising candidates.
@@
The insights gained from this study contribute to the development of high-performance Na-ion batteries, offering a viable and sustainable alternative to Li-ion technology for future energy storage applications. By leveraging machine learning-driven material discovery, this work paves the way for the rational design of next-generation battery materials, supporting advancements in clean energy technologies and grid-scale storage solutions.
Acknowledgment
@@
This work was supported by the Korea Institute of Science and Technology (Grant number 2E31851), GKP (Global Knowledge Platform, Grant number 2V6760) project of the Ministry of Science, ICT and Future Planning.
4 Conflicts of interest
The authors declare no conflict of interest. {suppinfo}
@@ 4 Conflicts of interest
The supporting information file includes the definition of R2 term, list of elemental features used for the building of the DNN model, computational methodology and expressions used for calculation of formation energy and voltage of new compositions. It also contains the structural information of two specific compositions, O3 and P2-type ordering (Figure S1). Formation energy of all the new compositions of O3 and P2-type ordering are presented in Figure S2.

### v3

@@ 4.4.1 Voltage comparison (ML vs DFT)
4.4.2 DFT & Boltzmann transport analysis
Density functional theory (DFT) calculations were employed to evaluate the electrochemical properties of the newly designed cathode materials: O3-type NaxMn0.61Ni0.28Cr0.11O2 and P2-type NaxMn0.5Fe0.28Cu0.22O2, where x = 1, 0.67 and 0.33. The corresponding voltage profiles during desodiation (i.e., discharge from x = 1 to x=0) are presented in Figure 11. The calculated voltage range for the O3-type material spans from 2.88 to 4.59 V, while the P2-type composition exhibits a slightly wider range from 2.75 to 4.97 V. The stepped features observed in the voltage profiles indicate the possibility of intermediate phase transitions during the desodiation process. The structural analysis (inset of Figure 11) reveals that the composition of type O3 undergoes a coordination transformation from octahedral (at x = 1) to mixed octahedral–prismatic (at x = 0.67) and finally to purely prismatic coordination at x=0.33, consistent with voltage steps and indicating progressive phase transformation.
In contrast, the P2-type NaxMn0.5Fe0.28Cu0.22O2 maintains a stable prismatic Na coordination throughout the desodiation process (from x = 1 to x=0.33), suggesting superior structural robustness and phase stability (inset of Figure 11) during cycling. This invariance in the coordination environment underscores its potential for high-performance and long-life Na-ion battery applications. The incorporation of Cu into the P2-type material has been reported to enhance electronic conductivity and introduce additional high-voltage redox activity via the Cu2+/Cu3+ couple 88, 89, 90, thereby supporting the observed broader voltage window. Importantly, the composition of the O3-type with intermediate Na content (i.e., Na0.67Mn0.61Ni0.28Cr0.11O2), which exhibits a coexistence of octahedral and prismatic Na coordination, may offer an advantageous balance between structural flexibility and electrochemical performance, making it a promising candidate for further optimization. Cr in the O3-type composition stabilizes high-voltage redox activity through the Cr3+/Cr4+ couple91, contributing to the increase in voltage beyond 4.0 V90. Overall, these DFT insights validate the viability of the ML-predicted materials and provide a strong foundation for future experimental studies and the design of next-generation Na-ion battery cathodes.
Figure 11: DFT calculated voltage profile of O3-NaxMn0.61Ni0.28Cr0.11O2 (black) and P2-NaxMn0.5Fe0.28Cu0.22O2 (red) during desodiation (x=1 to 0). Structural insets show coordination changes: O3-type undergoes a transition from octahedral to prismatic geometry, while P2-type retains prismatic coordination, indicating superior structural stability.
Figure 12: Spin-polarized total and projected density of states (DOS) for O3-type NaxMn0.61Ni0.28Cr0.11O2 at (a) x = 1, (b) x = 0.67, and (c) x = 0.33, and for P2-type NaxMn0.5Fe0.28Cu0.22O2 at (d) x = 1, (e) x = 0.67, and (f) x = 0.33. The evolution of the DOS with Na content reveals a narrowing band gap in the O3 system and metallic character in the P2 system, indicating improved electronic conductivity upon desodiation and the potential for high-rate capability in the P2-type material.
To gain deeper insight into the electronic structure and conductivity of the ML-designed cathode materials, we computed the total and projected density of states (DOS) for both O3- and P2-type compositions with different sodium contents (x = 1, 0.67, and 0.33), as shown in Figure 12. O3-type Na0.67Mn0.61Ni0.28Cr0.11O2 has higher DOS at Fermi level that can facilitate faster electron transport, as validated by the electrical conductivity data in Figure 13, can lead to improved power density and faster charging/discharging rates. For the O3-type NaxMn0.61Ni0.28Cr0.11O2 (Figures 12 a–c), a narrowing of the band gap is observed upon desodiation, particularly from x = 1 to x = 0.33, indicating enhanced electronic conductivity at higher states of charge. The contributions from Ni and Cr 3d states dominate near the Fermi level, suggesting their active participation in the redox process. In contrast, the P2-type NaxMn0.5Fe0.28Cu0.22O2 (Figures 12 d–f) exhibits a dramatic shift in the density of states (DOS) profile near the Fermi level, transitioning from metallic to semiconductor-like behavior. Notably, finite electronic states persist across all Na concentrations, with significant contributions at the Fermi level even at x = 1. Both Na0.67Mn0.5Fe0.28Cu0.22O2 and Na33Mn0.5Fe0.28Cu0.22O2 have large DOS near the Fermi level implying the availability of more number of states for electrons to occupy which reduces the energy barrier for electron excitation. The Cu and Fe 3d orbitals prominently contribute near the Fermi energy, indicating favorable electronic conductivity and robust charge-transport behavior during cycling. These electronic features support the observed voltage performance and suggest that the P2-type material may offer superior rate capability and electronic transport compared to the O3-type counterpart.
Figure 13: Electronic conductivity with respect to relaxation time ($\sigma/\tau$) as a function of chemical potential ($\mu$) at 300 K for (a) O3-type and (b) P2-type compositions.
We further evaluate the electronic transport characteristics of materials designed with ML within the constant relaxation time approximation using BoltzTraP code. The transport coefficient, electronic conductivity normalized by relaxation time ($\sigma/\tau$) as a function of chemical potential ($\mu$) was calculated at 300 K, as shown in Figure 13. The calculated electronic conductivity ($\sigma/\tau$) at 300 K for O3- and P2-type compositions reveals distinct trends across varying Na concentrations (x = 1, 0.67, and 0.33). For both structural types, $\sigma/\tau$ shows a non-monotonic dependence on the chemical potential ($\mu$), with a notable enhancement for intermediate Na content (x = 0.67). Specifically, O3-NaxMn0.61Ni0.28Cr0.11O2 exhibits the highest $\sigma/\tau$ near the Fermi level ($\mu\approx 0~\mathrm{eV}$), indicating favorable carrier mobility and improved electronic transport at this state of charge. Enhanced electrical conductivity leads to improved electrochemical behavior of the cathode 92. A similar trend is observed for P2-NaxMn0.5Fe0.28Cu0.22O2, albeit with a slightly lower magnitude. The enhanced $\sigma/\tau$ at x = 0.67 is correlated with mixed coordination environments and possible phase co-existence, which can facilitate a more effective charge delocalization. In contrast, compositions at x = 1 and 0.33 exhibit a reduced $\sigma/\tau$, likely due to more localized electronic states or increased structural distortion at high and low Na concentrations. These insights suggest that Na0.67 stoichiometry in both O3 and P2 frameworks may offer an optimal balance between electronic conductivity and structural integrity, contributing to superior rate capability and overall battery performance.
4.4.3 d-band center trends and electrochemical implications
To better understand the origin of the electrochemical behavior, we evaluated the d-band center ($\epsilon_{d}$) positions of both O3- and P2-type materials at different Na concentrations. The d-band center is defined via 93, 94,
$\varepsilon_{d}\;=\;\frac{\int_{-\infty}^{E_{F}}E\,D_{d}(E)\,dE}{\int_{-\infty}^{E_{F}}D_{d}(E)\,dE}$ (5)
Where $D_{d}(E)$ is the projected DOS onto transition-metal d states. $E_{F}$ is the Fermi energy.
An upward shift of the d-band center towards the Fermi energy level led to an enhanced Na+ desorption behavior in the KNVO / NVO@C heterostructure (caused by the electric field of the interface at the heterojunction), which consequently accelerated Na+ migration and facilitated diffusion of Na+95. We observe a similar upshift in the d-band center with desodiation in the compositions of the O3 and P2 types for x = 1 to x = 0.67. O3-type NaxMn0.61Ni0.28Cr0.11O2 exhibits a relatively shallow $\epsilon_{d}$, shifting from –1.305 eV at x = 1 to –1.246 eV at x = 0.33. In contrast, the P2-type NaxMn0.5Fe0.28Cu0.22O2 demonstrates deeper $\epsilon_{d}$ values, ranging from –1.862 eV to –1.965 eV over the same desodiation range. These trends are consistent with the theory of the d-band, where a higher (less negative) $\epsilon_{d}$ correlates with an increased hybridization between the transition metal (TM) d states and the O 2p orbitals, facilitating electronic conductivity and improving TM redox activity. For instance, O3-type materials, particularly at x = 0.67 ($\epsilon_{d}$ = –1.217 eV), present a favorable balance between redox potential and electronic transport, consistent with the enhanced $\sigma/\tau$ observed in Figure 13. This observation aligns with earlier studies [e.g. Ref96, Ref97], which link shallow d-band centers to improved electrode kinetics and higher average voltages in layered oxide cathodes.
Figure 14: $\varepsilon_{d}$ as a function of average cell voltage for O3–NaxMn0.61Ni0.28Cr0.11O2 and P2-NaxMn0.5Fe0.28Cu0.22O2
Finally, Fig.14 illustrates the evolution of the d-band center ($\varepsilon_{d}$) as a function of average cell voltage for the O3–NaxMn0.61Ni0.28Cr0.11O2 and P2-NaxMn0.5Fe0.28Cu0.22O2 cathode compositions. In both systems, $\varepsilon_{d}$ exhibits a non-monotonic dependence on sodium content, peaking near x=0.67, which corresponds to an intermediate state of desodiation. This shallowest $\varepsilon_{d}$ suggests enhanced hybridization between the transition-metal d states and oxygen 2p orbitals, facilitating improved electronic conductivity and redox activity. For the O3 phase, $\varepsilon_{d}$ remains consistently higher than in the P2 counterpart, indicating stronger TM–O covalency and potentially more delocalized charge transport pathways. Notably, the observed peak in $\varepsilon_{d}$ coincides with the computed maxima in electronic conductivity ($\sigma/\tau$) mentioned earlier, reinforcing the notion that the d-band center serves as a meaningful descriptor linking electronic structure with electrochemical performance.
5 Conclusions
@@
This study demonstrates the power of combining machine learning (ML) with first-principles simulations to accelerate the discovery and optimization of next-generation Na-ion battery cathodes. Our deep neural network (DNN) model, trained on Density Functional Theory (DFT) data, achieves a mean absolute error (MAE) of 0.24V, outperforming existing voltage prediction models with a high level of precision and a deviation of less than 5%. To validate the predictive capacity of our model, we compared its voltage estimations with DFT-calculated values for two newly investigated cathode compositions: O3-Na0.67Mn0.61Ni0.28Cr0.11O2 and P2-Na0.67Mn0.5Fe0.28Cu0.22O2. The excellent agreement between the ML-predicted and DFT-computed voltages confirms the reliability of our approach in capturing the electrochemical behavior of Na-ion materials. By significantly reducing the reliance on time-consuming and resource-intensive experimental methods, our approach provides a systematic framework for identifying promising cathode materials with optimized electrochemical properties. This data-driven methodology enables rapid screening of novel compositions, guiding experimental efforts toward the most promising candidates.
@@
The insights gained from this study contribute to the development of high-performance Na-ion batteries, offering a viable and sustainable alternative to Li-ion technology for future energy storage applications. Using machine learning-driven material discovery, we have helped pave the way for the rational design of next-generation battery materials, supporting advances in clean energy technologies and grid-scale storage solutions.
Acknowledgements
@@
This work was supported by the Korea Institute of Science and Technology (Grant number 2E31851), GKP (Global Knowledge Platform, Grant number 2V6760) project of the Ministry of Science, ICT and Future Planning. The authors thank Shubhayu Das of RWTH Aachen for useful discussions.
Conflicts of interest
@@
The authors declare that they have no conflict of interest. {suppinfo} The supporting information file includes the definition of R2 term, list of elemental characteristics used for the building of the DNN model. It also contains structural information on two specific compositions, the O3 and P2-type ordering (Figure S1). The formation energy of all the new O3 and P2-type ordering compositions is presented in Figure S2.